One strategy in a snowball fight is to throw a snowball at a high angle over level ground. While your opponent is watching the first one, a second sbowball is thrown at a low angle timed to arrive before or at the same time as the first one. Assume both snowballs are thrown with a speed 25.0 m/s. The first one is thrown at an angle of 70.0 degrees with respect to the horizontal.

A. At what angle should the second snowball be thrown to arrive at the same point as the first?
B. How many seconds later should the second snowball be thrown after the first to arrive at the same time?
1
Expert's answer
2018-09-24T12:25:08-0400
a.
θ=90-70=20.0°.
b.
t=2v/g sin⁡θ
t_1=(2(25))/9.8 sin⁡70=4.79 s.
t_2=(2(25))/9.8 sin⁡20=1.75 s.
Hence, the second snowball should be thrown after the first by
Δt=t_1-t_2=4.79-1.75=3.04 s.